OpenStudy (anonymous):

complete ordered pair (-2, ) so that it is a solution to the equation y = -x +4

OpenStudy (anonymous):

you know that x=-2, so substutite x=-2 into the original equation y= -x+4, to get y. that will get you the set of ordered pair

OpenStudy (anonymous):

does that mean I make the -2 now positive ?

OpenStudy (anonymous):

so you will have y= -(-2)+4=+2+4=6